How can Php 42,000 be invested part at 4% and rest at 7% to makr a yearly interest equal to 5% of Php 42,000
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 7.00% per annum ------------ Amount invested = y
42000
Interest----- 2100.00

Part I 4.00% per annum ---x
Part II 7.00% per annum ---y
Total investment
x + 1 y= 42000 -------------1
Interest on both investments
4.00% x + 7.00% y= 2100
Multiply by 100
4 x + 7 y= 210000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -168000.00
Add this to (2)
0 x 3 y= 42000
divide by 3
y = 14000
Part I 4.00% $ 28000
Part II 7.00% $ 14000

CHECK
28000 --------- 4.00% ------- 1120.00
14000 ------------- 7.00% ------- 980.00
Total -------------------- 2100.00

m.ananth@hotmail.ca